OpenStudy (anonymous):

solve 2x-3/x + 4 = x - 3/x

OpenStudy (anonymous):

\[\frac{2x-3}{x+4}=\frac{x-3}{x}\] start by cross multiplying to get \[x(2x-3)=(x+4)(x-3)\] and solve that one

OpenStudy (anonymous):

sorry i think u misinterpreted my equation. it is (2x-3) / (x) + (4) = (x) - 3/(x)

OpenStudy (anonymous):

i'm in pre calculus

OpenStudy (anonymous):

my directions are solve the equation

OpenStudy (anonymous):

perhaps \[\frac{2x-3}{x}+4=x-\frac{3}{x}\] like that?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

since \(x\) cannot be equal to zero, multiply both sides by \(x\) and start with \[2x-3+4x=x^2-3\] and solve that one

OpenStudy (anonymous):

6x = x^2

OpenStudy (anonymous):

what do i do from here?

OpenStudy (anonymous):

r u there?

OpenStudy (anonymous):

\[6x = x^2 \] \[x^2-6x=0\] \[x(x-6)=0\] so \(x=0\) which is impossible or \(x=6\)

OpenStudy (anonymous):

got it! THANKS